# Rethinking Causal Discovery Through the Lens of Exchangeability

### v1

@@ 5 SynthNN
Having created the synthetic dataset explained in Section 4.1, a new causal discovery method named SynthNN is proposed that simply consists of a neural network trained on the generated data.
In each example, the data consists of a table, where there is a connection between each pair, but there is no inherent order between samples. This is given by the defining properties of exchangeable data (and consequently of i.i.d. as well). Consequently, it is not clear at first how the input data should be fed into the neural network. Therefore, four different strategies were tested:
@@
    Building an image from the data and processing it using a convolutional neural network (CNN). This image corresponds exactly to the images shown when plotting the generated data. Out of the four, this is the only method that was applied to the problem of classifying cause-effect pairs by Singh et al. (Singh et al., 2017). Therefore, the designed network is generally based on this earlier implementation.
At first look, the second and fourth methods have a clear advantage: they are inherently invariant in the number of pairs, which is crucial to the classification problem at hand. Alternatively, the other two would require padding so that the dimension of all examples matches. However, after implementing the four different methods, the fourth, based on image processing, clearly stood out as the most promising one. Consequently, the others were dropped and the focus from now on will be on this one. The full details of the SynthNN are in Appendix C.
The key idea behind SynthNN is that if the synthetic dataset is representative of the Tübingen collection, by training a neural network on the former, it can achieve good results in the latter. However, these displayed significant variance based on the random weight initialization performed by the neural network. Consequently, the distribution of the obtained AUROC and accuracy in the training, validation, and Tübingen sets can be seen in Figure 3.
@@
The training and validation metrics displayed in Figure 3 show that the network is consistently learning the correct causal structure present in the synthetically generated examples. However, the performance obtained in the Tübingen data are far more varied, with AUROC having a standard deviation of $2.0\%$ and accuracy of $3.2\%$. This variation can be attributed to the weight initialization in the beginning of the training process. Nevertheless, if the training data were fully representative of the Tübingen dataset, then the random initialization would only serve to break the network symmetries and would always converge to the same result.
@@
Furthermore, Singh et al. (Singh et al., 2017) trained a very similar neural network directly on the Tübingen dataset, having achieved an AUROC of $76.9\%$ and an accuracy of $73.3\%$. Naturally, these results represent the extent to which a neural network (with this architecture) can properly learn the causal nature present in the Tübingen data. Consequently, it should represent an upper bound for the possible performance of any similar neural network trained on synthetic data. Therefore, since the results obtained by SynthNN are not far from those obtained by training directly on the Tübingen dataset, it is reasonable to conclude that most of the remaining error should be attributed to limitations of the neural network itself.
On the other hand, it also means the synthetic dataset should represent the dynamics of the Tübingen data accurately, at least as far as the neural network is able to capture the causal relationship between data. One can think of SynthNN as offering a lower bound on how representative the synthetic dataset is of the Tübingen data and thus hopefully of the real world.
Regardless of the point of view, it seems clear that the results obtained by SynthNN provide additional evidence of how representative the developed synthetic dataset is of real-world data. As such, this method serves as further validation not only of the dataset itself but also of its underlying assumption: the central role of exchangeability in causal discovery. This is even stronger, taking into account the many unintentional additional assumptions present in the design choices of both the synthetic dataset generation algorithm and the neural network architecture.
Furthermore, SynthNN is a causal discovery method in full right, since it estimates the causal direction of real-world data, only based on first principles. This is in stark contrast to the CNN from Singh et al. (Singh et al., 2017), which is itself trained on the Tübingen data. However, unlike other methods, such as LiNGAM or bQCD, it does not have such a direct downstream application to the problem of causal representation learning. While LiNGAM and bQCD’s assumptions have concise mathematical representations and thus can be more easily fit into a machine learning pipeline, SynthNN’s key assumption lies in the representational power of the synthetic dataset, which is harder to integrate. Nevertheless, SynthNN clearly shows the validity of developing a causal discovery method by training a neural network (or any other machine learning algorithm) on a dataset that is believed to represent real-world data well.
In conclusion, we believe SynthNN shows itself not only to be a valid method for causal discovery, but also further proves the central thesis of this paper: that i.i.d. causal discovery should be reframed as exchangeable.
@@ 6 Conclusion
We have shown that the traditional i.i.d. framework for causal discovery is most naturally and more powerfully understood as exchangeable. This was first substantiated through both conceptual and practical arguments (notably, the canonical real‑world benchmark, the Tübingen cause–effect pairs, aligns better with exchangeability than strict i.i.d.). Afterwards, to make these ideas more concrete, we introduced a novel synthetic dataset that enforces only exchangeability—eschewing stronger i.i.d. constraints—and showed that its statistical properties align more closely with the Tübingen data than any prior synthetic benchmark. Complementing the dataset, we presented a neural‑network–based causal‑discovery algorithm trained solely on our exchangeable examples. This model performs similarly to other relevant and current methods on real‑world data, demonstrating the viability of an exchangeability‑only approach. Even though the broader implications of adopting exchangeability in causal‑discovery practice remain open to be fully explored, this work opens the door to novel dependencies, structures, and methodologies that more accurately reflect the complexities of real‑world data.
@@ Appendix A Tübingen Statistical Assumptions Analysis
Table 4: Tübingen Analysis Data
@@
the solar radiation in W/m2 the daily average temperature of the air measured at the same location and the same days latitude timeseries
@@ Appendix B Synthetic Dataset Generation
B.1 Functions Specifications
Throughout this appendix, $x\in[0,1]$ denotes the min–max rescaled output produced by Algorithm 1. All latent parameters (collectively denoted $\psi$) are rescaled draws of a common latent variable sampled from one of the three base priors (Uniform, Normal, Rayleigh) (see Subsection 4.2). Additionally, since the latent variables were designed to interact with the input in complex ways, the common approach was to choose function hyperparameters that change the shape of the function. However, there still exists an option to add multiplicative noise, which will not be considered here. Lastly, in functions whose design is monotone for simplicity, a Bernoulli fair coin is tossed once per example to decide whether to flip the sign of $y$ (and, for some mechanisms, of $x$ as well) so that both monotone orientations occur with equal frequency. Now, in greater detail, the eight implemented causal functions are:
@@ B.1 Functions Specifications
    Exponential ($f_{\mbox{exp}}$): this implementation starts by sampling the latent variable rescaling parameters:
    $\mu_{\psi}\sim F_{\mu}(\mu_{\mu},\sigma_{\mu});\sigma_{\psi}\sim F_{\sigma}(\mu_{\sigma},\sigma_{\sigma})$ (10)
    where the choices for $\mu_{\mu},\sigma_{\mu},\mu_{\sigma},\sigma_{\sigma}$ and the functions $F_{\mu},F_{\sigma}$ constitute hyperparameters. Consequently, $y$ can be computed as
    $y_{i}=e^{x_{i}a_{i}},a=\mbox{Rescale}(\psi;\mu_{\psi},\sigma_{\psi}).$ (11)
@@
    Logarithmic ($f_{\log}$): the implementation of the logarithmic function closely resembles that of the exponential function above; the only change lies in the final function itself, which is
    $y_{i}=\log(x_{i}+a_{i}).$ (12)
@@
    Inversely proportional ($f_{\mbox{inv}}$): this function also mimics the exponential function in terms of obtaining the rescaled version of the latent variable. Its final expression is
    $y_{i}=\frac{1}{x_{i}+a_{i}}.$ (13)
@@
    Power law ($f_{\mathrm{pow}}$): the power law function is the last one that resembles the exponential in its architecture. Similarly, the expression used to compute $y$ from the rescaled $\psi$ distribution and $x$ is
    $y_{i}=x_{i}^{a_{i}}.$ (14)
@@
    Linear ($f_{\mbox{lin}}$): since $f$ must be strictly increasing in $\psi$ (in accordance with our design choice), in order to generate samples using linear functions it suffices to define the distribution at the two endpoints (0 and $1$ respectively)
    $\begin{gathered}\sigma_{0},\sigma_{1}\sim\mbox{Inv-Gamma}\bigl(\gamma_{m},\,\gamma_{v}\bigr)\\[6.0pt] \begin{aligned} \mu_{0}&=0\qquad&\mu_{1}&=\tan\phi\qquad&\phi&\sim\mathcal{U}(0,2\pi)\end{aligned}\end{gathered}$ (15)
    Consequently, the $\psi$ distribution is rescaled according to the sampled values above
    $a=\mbox{Rescale}(\psi;\mu_{0},\sigma_{0})\qquad b=\mbox{Rescale}(\psi;\mu_{1},\sigma_{1}),$ (16)
    where “Rescale” denotes the operation that adjusts the $\psi$ distribution to have mean $\mu$ and standard deviation $\sigma$, while preserving its original shape and point‐mass locations. Having the distributions at the two endpoints ($a$ and $b$), it is possible to obtain $y_{i}$
    $y_{i}=a_{i}+(b_{i}-a_{i})x_{i}.$ (17)
@@
    Piecewise linear ($f_{\mbox{mix}}$): the piecewise linear function can be implemented based on the linear function with two small adaptations:
    • There is an extra hyperparameter: $\mbox{max}_{K}$, which denotes the maximum number of slices the piecewise model should have; then, $K\sim\mathcal{U}(2,\mbox{max}_{K})$.
    • The smaller size of each slice (when compared to the entire $[0,1]$ interval) requires adapting the rescaling average $\mu$ to ensure the function is continuous:
      $\mu_{k+1}=\mu_{k}+\tan(\phi)(x_{k+1}-x_{k}),$ (18)
      where both $k$ and $k+1$ define the endpoints of each slice.
@@
    Brownian‐like motion ($f_{\mbox{brown}}$): the Brownian-like function can be seen has a piecewise linear function with two different adaptations:
    • Each slice only contains one point. Thus, there are $N_{S}$ slices.
    • In order to avoid constant oscillation, an extra momentum term is added to the computation of the rescaling variables to ensure the function is smooth and differentiable,
      $\mu_{k+1}\sim N\left(\mu_{k}+\frac{d\mu}{dt}\Delta_{t},\sqrt{\frac{\Delta_{t}^{P}}{P}}\right),$ (19)
      where $\frac{d\mu}{dt}$ is computed using a polynomial approximation, $P$ is a hyperparameter and the expression for the standard deviation is obtained from that of integrated Brownian motion.
@@
    Polynomial ($f_{\mbox{poly}}$): finally, the polynomial function starts by sampling the order of the polynomial: $o\sim\mathcal{U}(2,\mbox{max}_{o})$, where the maximum value ($\mbox{max}_{o}$) is an hyperparameter. Afterwards, $o+1$ random points are selected from $x$, which will represent our slices. Then, similarly to the piecewise linear function, we obtain the rescaled $\psi$ at each slice. Finally, for each $y_{i}$, given the corresponding resampled $\psi$,
    $\psi_{1_{i}},\psi_{k_{i}},\dots,\psi_{o_{i}}.$ (20)
    A polynomial of order $o$ is then fitted to the following pairs of points:
    $p_{i}=\mbox{fit}(x_{[k]},\psi_{[k]_{i}}).$ (21)
    Finally, $y_{i}$ can be computed as:
    $y_{i}=p_{i}(0)+p_{i}(1)x+p_{i}(2)x^{2}+...p_{i}(o)x^{o}.$ (22)

### v2

@@ 5 SynthNN
Having created the synthetic dataset, as explained in Section 4.1, a new causal discovery method named SynthNN is proposed, consisting of a neural network trained on the generated data. In each example, tabular data is provided, containing unordered causal pairs generated by Algorithm 1. The unordered nature of the pairs stems from the definition of exchangeability. Consequently, it is not clear at first how the input data should be fed into the neural network. Therefore, four different strategies were tested:
@@
    Building an image from the data and processing it using a convolutional neural network (CNN). This image corresponds exactly to the images shown when plotting the generated data. Out of the four, this is the only method that had been previously applied to the problem of classifying cause-effect pairs by Singh et al. (Singh et al., 2017). Therefore, the designed network is generally based on this earlier implementation.
At first look, the second and fourth methods have a clear advantage: they are inherently invariant in the number of pairs, which is crucial to the classification problem at hand. Alternatively, the other two would require padding so that the dimensions of all examples match. However, after implementing the four different methods, the fourth, based on image processing, clearly stood out as the most promising one. Consequently, the others were dropped, and the focus will be on this one. The full details of SynthNN, the implemented neural network, are in Appendix C.
The key idea behind SynthNN is that if the synthetic dataset is representative of the Tübingen collection, by training a neural network on the former, it can achieve good results in the latter. However, these results displayed significant variance based on the random weight initialization performed by the neural network. Consequently, the distribution of the obtained AUROC and accuracy in the training, validation, and Tübingen sets can be seen in Figure 4.
@@
Furthermore, Singh et al. (Singh et al., 2017) trained a very similar neural network directly on the Tübingen dataset, having achieved an AUROC of $76.9\%$ and an accuracy of $73.3\%$. These results represent the extent to which a neural network (with this architecture) can properly learn the causal nature present in the Tübingen data. Consequently, it should represent an upper bound for the possible performance of any similar neural network trained on synthetic data. Since the results obtained by SynthNN are not far from those obtained by training directly on the Tübingen dataset, it is reasonable to conclude that most of the remaining error should be attributed to limitations of the neural network itself.
On the other hand, it also means the synthetic dataset should represent the dynamics of the Tübingen data accurately, at least as far as the neural network is able to capture the causal relationship between data. One can think of SynthNN as offering a lower bound on how representative the synthetic dataset is of the Tübingen data and thus of real-world causal systems.
Regardless of the point of view, it seems clear that the results obtained by SynthNN provide additional evidence of how representative the developed synthetic dataset is of real-world data. As such, this method serves as further validation, not only of the dataset itself, but also of its underlying assumption: the central role of exchangeability in causal discovery. This is an even stronger result when taking into account the many unintentional additional assumptions present in the design choices of both the synthetic dataset generation algorithm and the neural network architecture.
Lastly, SynthNN is a causal discovery method in full right, since it estimates the causal direction of real-world data, only based on first principles. This is in stark contrast to the CNN from Singh et al. (Singh et al., 2017), which is itself trained on the Tübingen data. However, unlike other methods, such as LiNGAM or bQCD, it does not have such a direct downstream application to the problem of causal representation learning. While LiNGAM and bQCD’s assumptions have concise mathematical representations and thus can be more easily fit into a machine learning pipeline, SynthNN’s key assumption lies in the representational power of the synthetic dataset, which is harder to integrate. Nevertheless, SynthNN clearly shows the validity of developing a causal discovery method by training a neural network (or any other machine learning algorithm) on a dataset that is believed to represent real-world data.
@@ 6 Conclusion
We have shown that the traditional i.i.d. framework for causal discovery is most naturally and more powerfully understood as exchangeable. This was first substantiated with the proposal of a new hierarchical and exchangeable causal model that simultaneously adheres to the Causal de Finetti’s theorem, while enabling a deeper interpretation of the possibilities behind this approach. Specifically, the proposed model takes into consideration uncertainty regarding both the causal mechanism and the distribution of inner latent variables. Explicitly modelling these uncertainties is crucial to obtain a fairer description of the complexities of the real world. This argument is further supported by gathering inspiration from the neighbouring domain of causal inference, while an analysis of the Tübingen dataset heavily suggests a high number of unknown distributions of plausible latent variables.
Afterwards, to make these ideas more concrete, we introduced a novel synthetic dataset that enforces only exchangeability—eschewing stronger i.i.d. constraints—and showed that its statistical properties align more closely with the Tübingen dataset than any prior synthetic benchmark. Complementing the dataset, we presented a neural‑network–based causal‑discovery algorithm trained solely on our exchangeable examples. This model performs similarly to other relevant and current methods on real‑world data, demonstrating the viability of an exchangeability‑only approach. Even though the broader implications of adopting exchangeability in causal‑discovery practice remain open to be fully explored, this work opens the door to novel dependencies, structures, and methodologies that more accurately reflect the complexities of real‑world data.
7 Author’s Statement
Funding Information: This work was partially supported by Instituto de Telecomunicações (own funds).
Author Contribution: Conceptualization, TB and MB. Formal development, TB. Coding and experiments, TB. Writing original draft, TB. Revision and editing, TB and MF. Supervision, MF.
Conflict of Interest: None.
@@ Appendix A Tübingen Statistical Assumptions Analysis
This appendix provides a detailed analysis of the Tübingen dataset. Specifically, we analysed whether each example should be more accurately modelled using exchangeability, and the proposed hierarchical model defined in Equation (18). More so, we focused on the second condition and whether there were any relevant latent variables with unknown distributions that fit neatly into the description of exchangeable processes. Table LABEL:tab:Tübingen_data reports this analyses, alongside possible examples of these latent variables.
Table 4: Tübingen Dataset Analysis. The first two columns contain the cause and effect definitions (as in the source material). The latter two provide possible examples of latent variables with unknown distributions and whether the sequence should be considered exchangeable.
@@
[…]
@@ Appendix B Synthetic Dataset Generation
In this appendix, we first provide details on the 8 implemented functional mechanisms, controlled by the exchangeable latent variable $\psi_{2}$. Afterwards, we provide 32 random examples of causal pairs generated by the presented synthetic process..
B.1 Considered Functional Causal Mechanisms
In order to describe the 8 families of mathematical functions considered to represent the causal mechanism, we must first discuss the representation of their parameters. These are also defined by $\psi_{2}$, and they take different forms, depending on the specific family of functions under consideration. The controllable hyperparameters controlling the distributions from which these specific function parameters are sampled will be called $\lambda$. These were tuned to ensure that different parameter samples resulted in significant changes to the shape of the function, while ensuring both $X_{n}$ and $\gamma_{n}$ made significant contributions to the value of $Y_{n}$. This introduced randomness results in a more diverse and robust dataset.
For four of the implemented functions, the parameters will define the mean and standard deviation of $\gamma$. Remember that $\gamma$ had been min-max scaled to the interval $[0,1]$. After sampling the function parameters, $\gamma$ is rescaled using $\{\mu_{\gamma},\sigma_{\gamma}\}$, to ensure it has the desired mean and standard deviation. This rescaled distribution will be called $a$. Consequently, the hyperparameters define the distributions (and respective parameters) of both $\{\mu_{\gamma},\sigma_{\gamma}\}$. In all, this can be written as:
$\lambda$ $=\{F_{\mu},F_{\sigma},\mu_{\mu},\sigma_{\mu},\mu_{\sigma},\sigma_{\sigma}\},$ (26)
$\mu_{\psi_{2}}\sim$ $F_{\mu}(\mu_{\mu},\sigma_{\mu});\,\,\sigma_{\gamma}\sim F_{\psi_{2}}(\mu_{\sigma},\sigma_{\sigma}),$ (27)
$a$ $=\mbox{Rescale}(\gamma;\mu_{\psi_{2}},\sigma_{\psi_{2}}),$ (28)
where $F_{\mu},F_{\sigma}$ define the probability distributions (either Gaussian or Uniform), and $\mu_{\mu},\sigma_{\mu},\mu_{\sigma},\sigma_{\sigma}$ their parameters.
In specific, given $a$, these four functions are defined as:
@@ B.1 Considered Functional Causal Mechanisms
    Exponential ($f_{\mbox{exp}}$):
    $Y_{n}=e^{X_{n}a_{n}}.$ (29)
@@
    Logarithmic ($f_{\log}$):
    $y_{n}=\log(x_{n}+a_{n}).$ (30)
@@
    Inversely proportional ($f_{\mbox{inv}}$):
    $y_{n}=\frac{1}{x_{n}+a_{n}}.$ (31)
@@
    Power law ($f_{\mathrm{pow}}$):
    $y_{n}=x_{n}^{a_{n}}.$ (32)
For these mathematical families, the function parameters serve only to alter $\gamma$’s distribution. So, it would be fair to wonder why not include these parameters directly in $\gamma$’s distribution, defined by $\psi_{1}$. The current approach was preferred because 1) it allows for more flexibility, since the distribution of $\gamma$ can depend on $X_{n}$; and 2) it clears up the logic, leaving all detailed function parameters to $\psi_{2}$.
Having said this, for three more families of functions, the parameters will simultaneously control the rescaling process of $\gamma$ (as is done for the first four), and specific function criteria. The latter will be discussed for each function in time. Regarding the first, these four models will implement, even if in different ways, the same linear approach to rescale $\gamma$. The core idea is to sample (in accordance with $\lambda$) the mean and standard deviations ($\mu_{\text{start}},\sigma_{\text{start}},\mu_{\text{end}}$ and $\sigma_{\text{end}}$), associated with two (previously chosen) values of $X$ (called $X_{\text{start}}$ and $X_{\text{end}}$). The mean and standard deviation for $\gamma$ will depend on $X_{n}$, and it will be given by the weighted average of the mean and standard deviations at $X_{\text{start}}$ and $X_{\text{end}}$, in proportion to their distance. Mathematically:
$a_{\text{start}}=\mbox{Rescale}(\gamma;\mu_{\text{start}},\sigma_{\text{start}}),\,\,a_{\text{end}}=\mbox{Rescale}(\gamma;\mu_{\text{end}}$ $,\sigma_{\text{end}}),$ (33)
$a_{n}=\frac{a_{\text{end}_{n}}(X_{\text{end}}-X_{n})+a_{\text{start}_{n}}(X_{n}-X_{\text{start}})}{X_{\text{end}}-X_{\text{start}}},$ (34)
Additionally, these three families of functions will always consider a linear relationship between $a_{n}$ (the rescaled $\gamma_{n}$) and $X_{n}$,
$Y_{n}=a_{n}X_{n}.$ (35)
For the following three mathematical functions, the specific parameters that must be defined are: 1) which values will $X_{\text{start}}$ and $X_{\text{end}}$ take, and 2) how are $\mu_{\text{start}},\sigma_{\text{start}},\mu_{\text{end}}$ and $\sigma_{\text{end}}$ sampled. Additionally, some families of functions will break up the real axis into multiple intervals, each with their own $X_{\text{start}}$ and $X_{\text{end}}$. Since each boundary will necessarily serve as the right boundary of an interval and the left of the next one, we will use the notation $X_{k}$ to represent these points. Having said this, the three families of functions under consideration are:
@@
    Linear ($f_{\mbox{lin}}$):
    $X_{\text{start}}=0,\,X_{\text{end}}=1,$ (36)
    $\sigma_{\text{start}},\sigma_{\text{end}}\sim\text{Inv-Gamma}\bigl(\nu_{m},\,\nu_{v}\bigr),$ (37)
    $\mu_{\text{start}}=0\qquad\mu_{\text{end}}=\tan\phi\qquad\phi\sim\mathcal{U}(0,2\pi),$ (38)
    where $\mbox{Inv-Gamma}\bigl(\nu_{m},\,\nu_{v}\bigr)$ stands for the Inverse Gamma distribution with fixed hyperparameters $\lambda=\{\nu_{m},\,\nu_{v}\}$, and $\mathcal{U}$ is the uniform distribution.
@@
    Piecewise linear ($f_{\mbox{mix}}$):
    $K\sim\mathcal{U}(2,\mbox{max}_{K}),$ (39)
    $\{X_{k}\}=\{X_{l}\sim\mathcal{U}(0,1)\}_{K}\cup\{0,1\},$ (40)
    $\sigma_{k}\sim\text{Inv-Gamma}\bigl(\nu_{m},\,\nu_{v}\bigr),\,\,\phi\sim\mathcal{U}(0,2\pi),$ (41)
    $\mu_{0}=0,\mu_{k+1}=\mu_{k}+\tan(\phi)(X_{k+1}-X_{k}),$ (42)
    where $\lambda=\{\mbox{max}_{K},\nu_{m},\,\nu_{v}\}$ are the function’s hyperparameters.
@@
    Brownian‐like motion ($f_{\mbox{brown}}$):
    $\{X_{k}\}=X,$ (43)
    $\mu_{0}=0,\qquad\mu_{k+1}\sim\mathcal{N}\left(\mu_{k}+\frac{d\mu_{k}}{dt}\Delta_{t},\sqrt{\frac{\Delta_{t}^{3}}{3}}\right),$ (44)
    $\hat{\sigma}_{0}\sim\text{Inv-Gamma}\bigl(\nu_{m},\,\nu_{v}\bigr),\qquad\sigma_{0}=\hat{\sigma}_{0},\,\,$ (45)
    $\hat{\sigma}_{k+1}=\hat{\sigma}_{k}+\frac{d\hat{\sigma}_{k}}{dt}\Delta_{t},\qquad\sigma_{k+1}\sim\mathcal{N}\left(\hat{\sigma}_{k+1},\sqrt{\frac{\Delta_{t}^{3}}{3}}\right)$ (46)
    where $\lambda=\{\nu_{m},\,\nu_{v}\}$ are the function specific hyperparameters, and $\frac{d\mu_{k}}{dt},\frac{d\hat{\sigma}_{k}}{dt}$ are computed using a polynomial approximation.
Finally, the approach for the polynomial function shares many similarities with the previous three. It also starts by obtaining the rescaled $\gamma$ for different values of $X$ (called $a_{k}$). Then, for each $n$, it will fit a polynomial based on the pairs $\{X_{k},a_{k_{n}}\}$. Then, it will use the obtained coefficients to obtain $Y_{n}$. Mathematically,
@@
    Polynomial ($f_{\mbox{poly}}$):
    $K\sim\mathcal{U}(2,\mbox{max}_{K}),$ (47)
    $\{X_{k}\}=\{X_{l}\sim\mathcal{U}(0,1)\}_{K}\cup\{0,1\},$ (48)
    $\sigma_{k}\sim\text{Inv-Gamma}\bigl(\nu_{m},\,\nu_{v}\bigr),\,\,\phi\sim\mathcal{U}(0,2\pi),$ (49)
    $\mu_{0}=0,\mu_{k+1}=\mu_{k}+\tan(\phi)(X_{k+1}-X_{k}),$ (50)
    $\{p_{n}\}=\text{fit}(\{X_{k},a_{k_{n}}\}),$ (51)
    $Y_{n}=p_{n}(0)+p_{n}(1)X_{n}+p_{n}(2)X_{n}^{2}+...p_{n}(K)X_{n}^{K}.$ (52)
    where $\lambda=\{\nu_{m},\,\nu_{v}\}$ are the function specific hyperparameters, and the fit function solves a determined system.
Additionally, in functions whose design is monotone, a Bernoulli fair coin is tossed once per example to decide whether to flip the sign of $Y$ (and, for some mechanisms, of $X$ as well) so that both monotone orientations occur with equal frequency.